OpenStudy (anonymous):

What is approximate positive zero of the polynomial function f(x) = 2x^4 - 5x - 16?

OpenStudy (cwrw238):

f(2) = 2(16) - 10 - 16 = 6 f(1) = 2 - 5 - 16 = -19 thus the zero lies between 1 and 2

OpenStudy (anonymous):

Thank you!

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

Wait, the answer choices that are most congruous given your statement are {0.14, 1.19}, and {0.19, 2.14}

OpenStudy (anonymous):

Nevermind, my apologies! I was looking at the wrong question. The answer choices are 1.97, 1.89, and 1.75. How do I figure this out?

OpenStudy (cwrw238):

plug in these values into the function and see which one is closest to zero